What is the hydronium ion concentration of 1 g of acetic acid in 200 ml of water?


1
Expert's answer
2021-03-17T06:44:23-0400

The molarity of the solution is:

M = n / V = m / (Mr V)

where M - molarity, n - number of moles, m - mass, Mr - molecular weight, V - volume.

From here:

M = 1 g / (60 g/mol × 0.2 L) = 0.83 M

According to the dissociation equation:

Ka = [H3O+] [CH3COO-] / [CH3COOH]

As dissociation constant of acetic acid equals 1.7×10−5, the hydronium ion concentration is:

[H3O+] × [CH3COO-] = Ka × [CH3COOH]

[H3O+]2 = Ka × [CH3COOH]

[H3O+] = (Ka × [CH3COOH])1/2 = (1.7×10−5 × 0.83 M)1/2 = 0.00376 M


Answer: 0.00376 M